OpenStudy (anonymous):

Solve the equation for y: 4x + 3y = -12.

OpenStudy (jiteshmeghwal9):

subtaract 4x both sides & then divide 3 both sides

OpenStudy (anonymous):

4x + 3y = -12 Subtract 4x from both sides. 3y = 4x - 12 Divide both sides by 3. y = 4/3x - 4

hartnn (hartnn):

Hi, \(\huge \color{red}{\text{Welcome to Open Study}}\ddot\smile\)

OpenStudy (mayankdevnani):

3y=-12-4x y=-12-4x/3

OpenStudy (mayankdevnani):

can you sollve it @casswebsterr

OpenStudy (anonymous):

y= -4 - 4/3 x

OpenStudy (mayankdevnani):